OpenStudy (anonymous):

Convert the given form of the function into the two other forms. Vertex form -> y - 5 = -2(x+1)^2 I think standard would be -> y=-2x^2-4x+3 Can anyone explain to me how you get intercept form? Info you might need: Standard form = y = ax^2 + bx + c Vertex form = y = a(x-h)^2 Intercept form = y = a(x-r1)(x-r2)

OpenStudy (anonymous):

set y = 0 and solve for x you will get \[0-5=-2(x+1)^2\] \[(x+1)^2=\frac{5}{2}\] \[x+1=\pm\sqrt{\frac{5}{2}}=\pm\frac{\sqrt{10}}{2}\] \[x=-1-\frac{\sqrt{10}}{2},x=-1+\frac{\sqrt{10}}{2}\]

OpenStudy (anonymous):

so in factored form you will get \[y=2(x+1+\frac{\sqrt{10}}{2})(x+1-\frac{\sqrt{10}}{2})\]